OpenStudy (anonymous):

Find an exact value. tan 75°

OpenStudy (anonymous):

also, \[\cos 19\pi /12\]

OpenStudy (anonymous):

tan 75° = tan (30 + 45) = (tan 30 + tan 45) / (1 − tan 30 tan 45) = 2 + sqrt(3)

OpenStudy (anonymous):

thanks